OpenStudy (anonymous):

What is 11/12 - 5/12

jimthompson5910 (jim_thompson5910):

11/12 - 5/12 = (11 - 5)/12 11/12 - 5/12 = ??

OpenStudy (anonymous):

I got the answer allready